# Mixed Near-field and Far-field Target Localization for Low-altitude Economy

### v1

@@ Abstract
In this paper, we study efficient mixed near-field and far-field target localization methods for low-altitude economy, by capitalizing on extremely large-scale multiple-input multiple-output (XL-MIMO) communication systems. Compared with existing works, we address three new challenges in localization, arising from 1) half-wavelength antenna spacing constraint, 2) hybrid uniform planar array (UPA) architecture, and 3) incorrect mixed-field target classification for near-field targets. To address these issues, we propose a new three-step mixed-field localization method. First, we reconstruct the signals received at UPA antennas by judiciously designing analog combining matrices over time with minimum recovery errors, thus tackling the reduced-dimensional signal-space issue in hybrid arrays. Second, based on recovered signals, we devise a modified multiple signal classification (MUSIC) algorithm (catered to UPA architecture) to estimate 2D angular parameters of both far- and near-field targets. Due to half-wavelength inter-antenna spacing, there exist ambiguous angles when estimating true angles of targets. In the third step, we design an effective classification method to distinguish mixed-field targets, determine true angles of all targets, as well as estimate the ranges of near-field targets. In particular, angular ambiguity is resolved by showing an important fact that the three types of estimated angles (i.e., far-field, near-field, and ambiguous angles) exhibit significantly different patterns in the range-domain MUSIC spectrum. Furthermore, to characterize the estimation error lower-bound, we obtain a matrix closed-form Cramér-Rao bounds for mixed-field target localization. Finally, numerical results demonstrate the effectiveness of our proposed mixed-field localization method, which improves target-classification accuracy and achieves a lower root mean square error than various benchmark schemes.
Index Terms:
Integrated sensing and communications, extremely large-scale MIMO, mixed-field target localization, low-altitude economy.
@@ I Introduction
Integrated sensing and communications (ISAC) has been listed by International Telecommunication Union Radio communication Sector (ITU-R) as one of the six major usage scenarios in future sixth-generation (6G) wireless networks, to accommodate a wide range of emerging applications such as smart home [1], low-altitude economy (LAE) [2], autonomous driving [3]. Among various design paradigms, communication-centric designs has received upsurging research interests in both academia and industry, which effectively leverage existing wireless infrastructures (e.g., base stations (BSs), wireless backhaul) for enabling ubiquitous and high-resolution sensing. Following this paradigm, we study in this paper how to leverage extremely large-scale multiple-input multiple-output (MIMO) (XL-MIMO) systems to achieve mixed near-field and far-field target location for LAE [4]. In particular, we propose efficient methods to address a set of localization challenges arising from 1) half-wavelength antenna-spacing constraint, 2) hybrid uniform planar array (UPA) architecture, and 3) incorrect mixed-field target classification.
@@ I-A Related Works
Accurate and wide-range unmanned aerial vehicles (UAVs) localization lays important foundations for various tasks in LAE such as UAV shows, UAV cargo delivery, etc. To achieve this, one promising approach is by leveraging XL-MIMO systems to expand the sensing coverage and achieve high sensing resolution-and-accuracy in both near-field and far-field. Compared with conventional massive MIMO systems, XL-MIMO increases the number of antennas by another order-of-magnitude (e.g., from 256 to 1024), thus substantially enhancing system spectrum efficiency and spatial resolution. Moreover, the Rayleigh distance, which specifies the boundary between radiative near-field and far-field, is greatly increased. For example, when the system operates at 28 GHz frequency bands, the Rayleigh distance of an 1-meter (m) array is up to 187 m. As such, in future XL-MIMO ISAC systems, there may appear the case where some communication users/sensing targets are located in the near-field of the BS, while the others are in the far-field. This thus necessaries efficient designs for mixed-field ISAC in XL-MIMO systems.
@@
Specifically, for far-field localization, various algorithms have been proposed to estimate target angles capacitating on typical communication arrays. Taking digital arrays of half-wavelength inter-antenna spacing as an example, efficient subspace methods such as multiple signal classification (MUSIC) [11] and estimating signal parameters via rotational invariance techniques (ESPRIT) [12] have been developed to estimate target angle, which essentially exploit orthogonality between signal and noise subspaces to distinguish impinging sources. Further, for UPA systems, the one-dimensional (1D) MUSIC and ESPRIT algorithms were extended in [13] and [14] to joint estimate two-dimensional (2D) angular parameters (i.e., elevation and azimuth angles). To reduce the computational complexity arising from 2D exhaustive search in 2D MUSIC algorithms, the authors in [15] proposed to first exploit a 2D discrete Fourier transform (DFT) algorithm to estimate coarse 2D-angles and then refine angle estimation resolution by using angle rotation techniques [15].
@@
Next, for near-field localization, spherical wavefronts provide a new degree-of-freedom (DoF) in range estimation, hence allowing for joint estimation of both the target angle and range even in the narrow-band, yet without deploying multiple anchors with strict network synchronization. This can be achieved by designing 2D (angle-and-range domain) MUSIC algorithm for uniform linear array (ULA) systems and 3D (elevation/azimuth angles and range domain) for UPA systems, both applicable to half-wavelength antenna spacing. As 2D/3D exhaustive grid-search needs to be performed for target location, these algorithms incur prohibitively high computational complexity in practice. To address this issue, the authors in [16] assumed the inter-antenna spacing being no greater than one-quarter wavelength, hence avoiding intricate angular ambiguity. Based on this, they proposed a reduced-dimensional MUSIC (RD-MUSIC) algorithm, which only requires 1D MUSIC spectrum search in the angular domain via splitting array steering matrix in terms of angle and range. Besides, by extracting anti-diagonal elements of the original covariance matrix, an effective modified MUSIC algorithm was proposed in [17] (yet still based on one-quarter wavelength antenna spacing), which decomposed 2D (angle-and-range domain) peak search into two 1D grid-search in the angular and range domains, respectively. Moreover, for UPA systems, the authors in [18] leveraged the anti-diagonal elements of signal covariance matrix to decouple the 3D MUSIC algorithm into 2D angle and 1D range estimation problems, whose complexity was further reduced by exploiting 2D DFT algorithm and angle rotation technique in [15].
@@
Last, for mixed-field localization, due to the coupling between received spherical and planar wavefronts, it is necessary to differentiate far-field and near-field targets. To achieve this, for ULA systems, the authors in [19, 17] proposed to firstly identify the directions-of-arrival (DoAs) of far-field sources by finding the peak of the 2D (angle-and-range) MUSIC spectrum in the angular domain via setting the range parameter $r\to\infty$ [19, 17]. Then, far-field signals are removed and the locations of near-field users are estimated from residual signals by using typical methods as for near-field localization. The authors in [20] addressed the mixed near-field and far-field target localization problem by a two-stage MUSIC algorithm, which, however, incurs high computational complexity. Furthermore, an enhanced approach was proposed in [21] by applying a matrix auto-differencing operation to the fourth-order cumulant (FOC) matrix for achieving superior classification and localization performance in such mixed scenarios. In view of the above works on mixed-field localization, there still exist several limitations as listed below.
@@
  • (Half-wavelength inter-antenna spacing) First, for low-complexity mixed-localization design, the existing works rely on a strong assumption that the inter-antenna spacing is no greater than one-quarter wavelength to avoid angular ambiguity, which, however, is incompatible to legacy wireless systems with half-wavelength spacing.
@@
  • (Hybrid UPA architecture) Second, existing angle/location estimation methods mostly considered digital array systems. Nevertheless, these methods may not be applicable to hybrid UPA systems (widely deployed in high-frequency wireless communication and/or ISAC systems), due to the reduced sensing DoFs after analog combining as well as more complex signal processing techniques for UPAs.
  • (Incorrect mixed-field target classification) The mixed-field target classification methods in [19, 17] may make false detection of near-field targets, when their ranges are relatively large (but still smaller than Rayleigh distance). This is because in this case, the near-field channel is highly correlated with far-field channel at the same angle (to be shown in Section III-C), thus making it possible to be wrongly detected as far-field target. Such target classification error will inevitably incur high target sensing/location errors.
@@ I-B Contribution and Organization
To address the above issues, in this paper, we consider an XL-MIMO target localization system for LAE as shown in Fig. 1, where the BS is equipped with a hybrid UPA array with half-wavelength inter-antenna spacing. Our target is to employ the XL-MIMO BS to localize both near-field and far-field UAVs in low altitude, where the UAVs actively send probing signals to facilitate localization at the BS. The main contributions of this paper are summarized as follows.
First, we propose a new and efficient framework to localize both near-field and far-field UAVs, which comprises three main steps. In Step 1, we judiciously design the analog combining matrices at the BS over time to recover the received signals at the UPA with minimum recovery errors, thus addressing the issue in the reduced-dimensional received signals in hybrid arrays. Then, in Step 2, we decouple the joint angle-and-range estimation into two sequential angle and range estimations, by extracting anti-diagonal elements of the original covariance matrix. A modified MUSIC algorithm is further devised to estimate candidate 2D-angular parameters for all targets, including the true far-field and near-field 2D angles as well as their corresponding ambiguous angles arising from half-wavelength inter-antenna spacing. In Step 3, we design an effective classification method to distinguish mixed-field targets, determine true angles of all targets, as well as estimate the ranges of near-field targets. In particular, we resolve angular ambiguity by showing an important fact that the three types of estimated angles have significantly different patterns in the range-domain MUSIC spectrum. In addition, to evaluate the lower-bound of the proposed algorithm, we characterize the Cramér-Rao Bounds (CRB) for mixed-field target localization. Moreover, extensive numerical results are presented to demonstrate the effectiveness of our proposed mixed-field target localization method. It is shown that the proposed algorithm can achieve nearly the same performance with the 3D MUSIC algorithm, while significantly reducing its computational complexity. Moreover, due to larger array aperture, the proposed mixed-field localization method achieves lower root mean square error (RMSE) than UPAs with one-quarter wavelength inter-antenna spacing. Besides, our proposed method demonstrates superior effectiveness in the mixed-field target classification.
[…]
@@ II System Model
We consider an XL-MIMO wireless sensing system shown in Fig. 1, where a BS equipped with a UPA is employed for localizing both near-field and far-field UAVs in low-altitude. The BS is placed in the $xy$-plane and centered at the origin, which consists of $N=N_{x}N_{y}$ antennas with $N_{x}$ and $N_{y}$ denoting the number of antennas along the $x$-axis and $y$-axis, respectively. Different from existing literature for mixed-field target localization assuming fully-digital architectures, we consider typical wireless communication infrastructures in this paper. First, the inter-antenna spacing of UPA along the $x$-axis and $y$-axis is half-wavelength, which are respectively denoted by $d_{x}=d_{y}=d_{0}=\frac{\lambda}{2}$, with $\lambda$ denoting the carrier wavelength. Second, the BS employs a hybrid array architecture, for which $N$ antennas are connected to $N_{\rm RF}\ll N$ radio frequency (RF) chains with each one connecting $U\triangleq{N}/{N_{\rm RF}}=M_{x}M_{y}$ phase shifters as illustrated in Fig. 1 (referred to as sub-connected architecture). Here, $U_{x}$ and $U_{y}$ are the number of PSs (assumed to be integers) along the $x$-axis and $y$-axis, respectively, and we assume $U={N}/{N_{\rm RF}}$ being an integer without loss of generality. Moreover, we consider the scenario where both near-field and far-field UAVs actively transmit uncorrelated signals to facilitate precise localization at the BS, which typically occurs in LAE, such as BS-assisted localization for UAV shows and cargo delivery, while the extension to the passive-sensing case will be discussed in Section III-D.
@@ II-A Channel Models
For the BS, the $n$-th antenna is located at $(n_{x}d_{0},n_{y}d_{0},0)$, where $n_{x}\in\mathcal{N}_{x}\triangleq\{-\frac{N_{x}-1}{2},\cdots,\frac{N_{x}-1}{2}\}$ and $n_{y}\in\mathcal{N}_{y}\triangleq\{-\frac{N_{y}-1}{2},\cdots,\frac{N_{y}-1}{2}\}$ denote the indices of antennas along the $x$-axis and $y$-axis, respectively. As such, the antenna index $n$ can be represented as $n=n_{x}+n_{y}N_{x}+\frac{N+1}{2}\in\mathcal{N}\triangleq\{1,2,\cdots,N\}$, and the Rayleigh distance of the UPA is given by $Z_{\rm R}=2(N_{x}^{2}d_{0}^{2}+N_{y}^{2}d_{0}^{2})/{\lambda}$. Let $K$ denote the number of UAVs in low-altitude, consisting of $K_{1}$ UAVs in the far-field and $K_{2}$ UAVs in the near-field of the UPA, whose channels are modeled as follows.
Far-field channel model: For far-field UAV $k\in\mathcal{K}_{1}\triangleq\{1,2,\cdots,K_{1}\}$, its range with the BS is larger than the Rayleigh distance $Z_{\rm R}$. For target localization in high-frequency bands, we mainly focus on the line-of-sight (LoS) channel path as the NLoS channel paths exhibit negligible power due to severe path-loss and shadowing [7]. Hence, the far-field channel model from UAV $k$ to the BS, denoted by $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}$, is given by
@@
where $h_{k}=\frac{\sqrt{\beta_{0}}}{r_{k}}$ denotes the complex channel gain with $\beta_{0}$ representing the reference channel power gain at a range of 1 m. Moreover, $\mathbf{a}(\theta_{{k}},\phi_{{k}})$ denotes the far-field steering vector with $\theta_{{k}}\in[0,\frac{\pi}{2}]$ and $\phi_{{k}}\in[-\pi,\pi]$ denoting the elevation and azimuth angles of the $k$-th UAV ($k\in\mathcal{K}_{1}$), respectively. In particular, $\mathbf{a}(\theta_{{k}},\phi_{{k}})$ is given by
@@
where $\mathbf{a}_{x}(\alpha_{{k}})\in\mathbb{C}^{N_{x}\times 1}$ and $\mathbf{a}_{y}(\beta_{{k}})\in\mathbb{C}^{N_{y}\times 1}$ represent the channel response vectors for the antennas in the $x$-axis and $y$-axis, respectively. Mathematically, we have $[\mathbf{a}_{x}(\alpha_{{k}})]_{n_{x}+\frac{N_{x}+1}{2}}=e^{\jmath\frac{2\pi}{\lambda}d_{0}n_{x}\alpha_{k}}$ and $[\mathbf{a}_{y}(\beta_{{k}})]_{n_{y}+\frac{N_{y}+1}{2}}=e^{\jmath\frac{2\pi}{\lambda}d_{0}n_{y}\beta_{k}}$, where parameters $\alpha_{k}$ and $\beta_{k}$ are defined as
$\alpha_{k}\triangleq\sin\theta_{{k}}\cos\phi_{{k}},\leavevmode\nobreak\ \leavevmode\nobreak\ \beta_{k}\triangleq\sin\theta_{{k}}\sin\phi_{{k}}.$
@@
Near-field channel model: For near-field UAV $k\in\mathcal{K}_{2}\triangleq\{K_{1}+1,K_{1}+2,\cdots,K\}$, its range $r_{k}$ is smaller than the Rayleigh distance $Z_{\rm R}$. As such, the near-field LoS channel from UAV $k\in\mathcal{K}_{2}$ to the BS can be modeled as [22]
@@
where $r_{k,n}$ denotes the range from the near-field UAV $k\in\mathcal{K}_{2}$ to the $n$-th antenna. Mathematically, $r_{k,n}$ can be expressed as
@@
for $k\in\mathcal{K}_{2}$, where $r_{k}$ represents the range between the UAV and the center of the BS. Further, by means of two-dimensional Fresnel approximation, $r_{k,n}$ in (5) can be approximated as [23]
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \!+\!\frac{n_{x}^{2}d_{0}^{2}(1-\alpha_{k}^{2})\!+\!n_{y}^{2}d_{0}^{2}(1-\beta_{k}^{2})}{2r_{k}}\!-\!\frac{n_{x}n_{y}d_{0}^{2}\alpha_{k}\beta_{k}}{r_{k}}.$
Then, the near-field channel steering vector for UAV $k\in\mathcal{K}_{2}$ can be rewritten as
@@
where $u_{n}(\alpha_{k},\beta_{k})\triangleq n_{x}d_{0}\alpha_{k}+n_{y}d_{0}\beta_{k}$ contains the angle information only, and $v_{n}(r_{k},\alpha_{k},\beta_{k})\triangleq\frac{n_{x}n_{y}d_{0}^{2}\alpha_{k}\beta_{k}}{r_{k}}-\frac{n_{x}^{2}d_{0}^{2}(1-\alpha_{k}^{2})+n_{y}^{2}d_{0}^{2}(1-\beta_{k}^{2})}{2r_{k}}$ is jointly determined by both the UAV angles and range.

### v2

# Mixed Near-field and Far-field Localization in Extremely Large-scale MIMO Systems

@@ Abstract
In this paper, we study efficient mixed near-field and far-field target localization methods in extremely large-scale multiple-input multiple-output (XL-MIMO) systems. Compared with existing works, we address two new challenges in target localization of MIMO communication systems via using decoupled subspace methods, arising from the half-wavelength antenna spacing constraint and hybrid uniform planar array (UPA) architectures. To this end, we propose a new three-step mixed-field localization method. First, we reconstruct the equivalent signals received at UPA antennas by judiciously designing analog combining matrices over time with minimum recovery errors. Second, based on recovered signals, we extend the modified multiple signal classification (MUSIC) algorithm to the UPA architectures by constructing a new covariance matrix of a virtual sparse UPA (S-UPA) to decouple the 2D angles and range estimation. Due to the structure of the S-UPA, there exist ambiguous angles when estimating true angles of targets. In the third step, we design an effective classification method to distinguish mixed-field targets, determine true angles of all targets, as well as estimate the ranges of near-field targets. In particular, angular ambiguity is resolved by showing an important fact that the three types of estimated angles (i.e., far-field, near-field, and ambiguous angles) exhibit significantly different patterns in the range-domain MUSIC spectrum. Furthermore, to characterize the estimation error lower-bound, we obtain a matrix closed-form Cramér-Rao bounds for mixed-field target localization. Finally, numerical results demonstrate the effectiveness of our proposed mixed-field localization method, which improves target-classification accuracy and achieves a lower root mean square error than various benchmark schemes.
@@ I Introduction
Sensing functionality at the base station (BS) has been listed by International Telecommunication Union Radio communication Sector (ITU-R) as one of the six major usage scenarios in future sixth-generation (6G) wireless networks, to accommodate a wide range of emerging applications such as smart home, low-altitude economy (LAE), autonomous driving [23, 22]. Among various design paradigms, communication-centric designs has received upsurging research interests in both academia and industry, which effectively leverage existing wireless infrastructures (e.g., BSs, wireless backhaul) for enabling ubiquitous and high-resolution sensing. Following this paradigm, we study in this paper how to leverage extremely large-scale multiple-input multiple-output (MIMO) (XL-MIMO) systems to achieve mixed near-field and far-field target localization [11]. In particular, we propose efficient methods to address a set of localization challenges by using decoupled subspace methods encountered in XL-MIMO communication systems, which arise from the half-wavelength antenna spacing constraint and the hybrid uniform planar array (UPA) architecture.
@@ I-A Related Works
Accurate and wide-range localization lays important foundations for various tasks in cellar networks such as beam management, user tracking, etc. To achieve this, one promising approach is by leveraging XL-MIMO systems to expand the sensing coverage and achieve high sensing resolution-and-accuracy in both near-field and far-field. Compared with conventional massive MIMO systems, XL-MIMO increases the number of antennas by another order-of-magnitude (e.g., from 256 to 1024), thus substantially enhancing system spectrum efficiency and spatial resolution. Moreover, the Rayleigh distance, which specifies the boundary between radiative near-field and far-field, is greatly increased. For example, when the system operates at $28$ GHz frequency bands, the Rayleigh distance of an 1-meter (m) array is up to 187 m. As such, in future XL-MIMO systems, there may appear the case where some communication users/sensing targets are located in the near-field of the BS, while the others are in the far-field. This thus necessaries efficient designs for mixed-field ISAC in XL-MIMO systems.
@@
Specifically, for far-field localization, various algorithms have been proposed to estimate target angles using typical communication arrays. Taking digital arrays of half-wavelength inter-antenna spacing as an example, efficient subspace methods such as multiple signal classification (MUSIC) [15] and estimating signal parameters via rotational invariance techniques (ESPRIT) [14] have been developed to estimate target angle, which essentially exploit orthogonality between signal and noise subspaces to distinguish impinging sources. Further, for UPA systems, the one-dimensional (1D) MUSIC and ESPRIT algorithms were extended in [1] and [18] to joint estimate two-dimensional (2D) angular parameters (i.e., elevation and azimuth angles). To reduce the computational complexity arising from 2D exhaustive search in 2D MUSIC algorithms, the authors in [4] proposed to first exploit a 2D discrete Fourier transform (DFT) algorithm to estimate coarse 2D angles and then refine angle estimation resolution by using angle rotation techniques [4].
@@
Next, for near-field localization, spherical wavefronts provide a new degree-of-freedom (DoF) in range estimation, hence allowing for joint estimation of both the target angle and range even in the narrow-band, yet without deploying multiple anchors with strict network synchronization [3]. This can be achieved by designing 2D (angle-and-range domain) MUSIC algorithm for uniform linear array (ULA) systems and 3D (elevation/azimuth angles and range domain) for UPA systems. As 2D/3D exhaustive grid-search needs to be performed for target location, these algorithms incur prohibitively high computational complexity in practice. To address this issue, the authors in [24] assumed that the inter-antenna spacing is no larger than one-quarter wavelength, hence avoiding intricate angular ambiguity. Based on this, they proposed a reduced-dimensional MUSIC (RD-MUSIC) algorithm, which only requires 1D MUSIC spectrum search in the angular domain via splitting array steering matrix in terms of angle and range. Besides, by extracting anti-diagonal elements of the original covariance matrix, an effective modified MUSIC algorithm was proposed in [6] (yet still based on one-quarter wavelength antenna spacing), which decomposed 2D (angle-and-range domain) peak search into two 1D grid-search in the angular and range domains, respectively. Moreover, for UPA systems, the authors in [8] leveraged the anti-diagonal elements of signal covariance matrix to decouple the 3D MUSIC algorithm into 2D angle and 1D range estimation problems, whose complexity was further reduced by exploiting 2D DFT algorithm and angle rotation technique in [4].
@@
Last, for mixed-field localization, due to the coupling between received spherical and planar wavefronts, it is necessary to differentiate far-field and near-field targets. To achieve this, for ULA systems, the authors in [10, 6] proposed to firstly identify the angle of far-field sources by finding the peak of the 2D (angle-and-range) MUSIC spectrum in the angular domain via setting the range parameter $r\to\infty$ [10, 6]. Then, far-field signals were removed and the locations of near-field users were estimated from residual signals by using typical methods as for near-field localization. The authors in [9] addressed the mixed near-field and far-field target localization problem by a two-stage MUSIC algorithm, which, however, incurs high computational complexity. Furthermore, an enhanced approach was proposed in [7] by applying a matrix auto-differencing operation to the fourth-order cumulant (FOC) matrix for achieving superior classification and localization performance in such mixed scenarios. In view of the above works on mixed-field localization, there still exist several limitations as listed below.
@@
  • (Half-wavelength inter-antenna spacing) For low-complexity mixed-localization design using subspace methods, the existing works rely on a strong assumption that the inter-antenna spacing is no larger than one-quarter wavelength to avoid angular ambiguity, which, however, is incompatible to legacy wireless systems with typically half-wavelength spacing. Moreover, larger number of ambiguous angles arising from half-wavelength inter constraints also further complicate the differentiation between far-field and near-field targets.
@@
  • (Hybrid UPA architecture) Existing angle/location estimation methods mostly considered fully digital architectures. Nevertheless, these methods may not be applicable to hybrid UPA systems (widely deployed in high-frequency wireless communication and/or ISAC systems), due to the reduced sensing DoFs after analog combining as well as more complex signal processing techniques for UPAs.
@@ I-B Contribution and Organization
To address the above issues, in this paper, we consider an XL-MIMO target localization system for mixed-filed targets as shown in Fig. 1, where the BS is equipped with a hybrid UPA array with half-wavelength inter-antenna spacing. Our target is to employ the XL-MIMO BS to localize both near-field and far-field targets, where the targets actively send probing signals to facilitate localization at the BS. The main contributions of this paper are summarized as follows.
First, we propose a new and efficient framework to localize both near-field and far-field targets, which comprises three main steps. In Step 1, we judiciously design the analog combining matrices at the BS over time to recover the equivalent signals (like a fully-digital architecture) received at the UPA with minimum recovery errors. Then, in Step 2, we extend the reduced-dimensional MUSIC algorithm to the UPA scenario by constructing an equivalent far-field covariance matrix of a virtual sparse UPA (S-UPA) from the original covariance matrix, which decouples the joint 3D angle-and-range estimation into sequential angle and range estimations. Subsequently, a decoupled 2D angular MUSIC algorithm is developed to further separate the estimation of elevation and azimuth angles, obtaining candidate 2D angles for all targets, including both true far-field and near-field angles as well as the corresponding ambiguous angles induced by the virtual S-ULA. In Step 3, we design an effective classification method to distinguish mixed-field targets, determine true angles of all targets, as well as estimate the ranges of near-field targets. In particular, we resolve angular ambiguity by showing an important fact that the three types of estimated angles have significantly different patterns in the range-domain MUSIC spectrum under the half-wavelength inter-antenna spacing constraint. In addition, to evaluate the lower-bound of the proposed algorithm, we characterize the Cramér-Rao Bounds (CRB) for mixed-field target localization. Moreover, extensive numerical results are presented to demonstrate the effectiveness of our proposed mixed-field target localization method. It is shown that the proposed algorithm can achieve nearly the same performance with the 3D MUSIC algorithm, while significantly reducing its computational complexity. In addition, due to larger array aperture, the proposed mixed-field localization method achieves a lower root mean square error (RMSE) than UPAs with one-quarter wavelength inter-antenna spacing.
Notations: Lowercase and uppercase boldface letters are used to represent vectors and matrices, such as $\mathbf{a}$ and $\mathbf{A}$. For vectors and matrices, the symbol $(\cdot)^{H}$ indicates the conjugate transpose operation, while the symbol $\mathbf{A}]_{i,j}$ and $\mathbf{a}]_{n}$ denote the $(i,j)$-th entry of the matrix $\mathbf{A}$ and the $n$-th entry of the vector $\mathbf{a}$, respectively. Additionally, calligraphic letters are employed to denote sets, while the symbol $\mathbb{Z}$ represents the integer set. The notations $\left|\cdot\right|$ and $\left\lVert\cdot\right\lVert$ refer to the absolute value of a scalar and the $\ell_{2}$ norm, respectively. The symbol $\mathbf{I}_{K}$ represents a $K$-dimensional identity matrix, while the symbol $\text{angle}(x)$ denotes the phase of the complex value $x$. Finally, the notation $\mathcal{CN}(\mu,\sigma^{2})$ represents the complex Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$
@@ II System Model
We consider an XL-MIMO wireless sensing system shown in Fig. 1, where a BS equipped with a UPA is employed for localizing both near-field and far-field targets. The BS is placed in the $xy$-plane and centered at the origin, which consists of $N=N_{x}N_{y}$ antennas with $N_{x}$ and $N_{y}$ denoting the number of antennas along the $x$-axis and $y$-axis, respectively. Different from existing literature for mixed-field target localization assuming fully-digital architectures, we consider typical wireless communication infrastructures in this paper. First, the inter-antenna spacing of UPA along the $x$-axis and $y$-axis is half-wavelength, which are respectively denoted by $d_{x}=d_{y}=d_{0}=\frac{\lambda}{2}$, with $\lambda$ denoting the carrier wavelength. Second, the BS employs a hybrid array architecture, for which $N$ antennas are connected to $N_{\rm RF}\ll N$ radio frequency (RF) chains with each one connecting $U\triangleq{N}/{N_{\rm RF}}=M_{x}M_{y}$ phase shifters as illustrated in Fig. 1 (referred to as sub-connected architecture). Here, $U_{x}$ and $U_{y}$ are the number of PSs (assumed to be integers) along the $x$-axis and $y$-axis, respectively, and we assume $U={N}/{N_{\rm RF}}$ being an integer without loss of generality. Moreover, we consider the passive sensing scenario (i.e., the BS does not actively transmit probing signals) where both near-field and far-field targets transmit uncorrelated signals to facilitate precise localization at the BS, while the extension to BS active sensing case will be discussed in Section III-D.
@@ II-A Channel Models
For the BS, the $n$-th antenna is located at $(n_{x}d_{0},n_{y}d_{0},0)$, where $n_{x}\in\mathcal{N}_{x}\triangleq\{-\frac{N_{x}-1}{2},\cdots,\frac{N_{x}-1}{2}\}$ and $n_{y}\in\mathcal{N}_{y}\triangleq\{-\frac{N_{y}-1}{2},\cdots,\frac{N_{y}-1}{2}\}$ denote the indices of antennas along the $x$-axis and $y$-axis, respectively. As such, the antenna index $n$ can be represented as $n=n_{x}+n_{y}N_{x}+\frac{N+1}{2}\in\mathcal{N}\triangleq\{1,2,\cdots,N\}$, and the Rayleigh distance of the UPA is given by $Z_{\rm R}=2(N_{x}^{2}d_{0}^{2}+N_{y}^{2}d_{0}^{2})/{\lambda}$. Let $K$ denote the number of targets, consisting of $K_{1}$ targets in the far-field and $K_{2}$ targets in the near-field of the UPA, whose channels are modeled as follows.
Far-field channel model: For far-field target $k\in\mathcal{K}_{1}\triangleq\{1,2,\cdots,K_{1}\}$, its range with the BS is larger than the Rayleigh distance $Z_{\rm R}$. For target localization in high-frequency bands, we mainly focus on the line-of-sight (LoS) channel path as the NLoS channel paths exhibit negligible power due to severe path-loss and shadowing [21]. Hence, the far-field channel model from target $k$ to the BS, denoted by $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}$, is given by
@@
where $h_{k}=\frac{\sqrt{\beta_{0}}}{r_{k}}$ denotes the complex channel gain with $\beta_{0}$ representing the reference channel power gain at a range of 1 m. Moreover, $\mathbf{a}(\theta_{{k}},\phi_{{k}})$ denotes the far-field steering vector with $\theta_{{k}}\in[0,\frac{\pi}{2}]$ and $\phi_{{k}}\in[-\pi,\pi]$ denoting the elevation and azimuth angles of the $k$-th target ($k\in\mathcal{K}_{1}$), respectively. In particular, $\mathbf{a}(\theta_{{k}},\phi_{{k}})$ is given by
@@
where $\mathbf{a}_{x}(\alpha_{{k}})\in\mathbb{C}^{N_{x}\times 1}$ and $\mathbf{a}_{y}(\beta_{{k}})\in\mathbb{C}^{N_{y}\times 1}$ represent the channel response vectors for the antennas in the $x$-axis and $y$-axis, respectively. Mathematically, we have $\mathbf{a}_{x}(\alpha_{{k}})]_{n_{x}+\frac{N_{x}+1}{2}}=e^{\jmath\frac{2\pi}{\lambda}d_{0}n_{x}\alpha_{k}}$ and $\mathbf{a}_{y}(\beta_{{k}})]_{n_{y}+\frac{N_{y}+1}{2}}=e^{\jmath\frac{2\pi}{\lambda}d_{0}n_{y}\beta_{k}}$, where parameters $\alpha_{k}$ and $\beta_{k}$ are defined as
$\alpha_{k}\triangleq\sin\theta_{{k}}\cos\phi_{{k}},\penalty 10000\ \penalty 10000\ \beta_{k}\triangleq\sin\theta_{{k}}\sin\phi_{{k}}.$
@@
Near-field channel model: For near-field target $k\in\mathcal{K}_{2}\triangleq\{K_{1}+1,K_{1}+2,\cdots,K\}$, its range $r_{k}$ is smaller than the Rayleigh distance $Z_{\rm R}$. As such, the near-field LoS channel from target $k\in\mathcal{K}_{2}$ to the BS can be modeled as [25]
@@
where $r_{k,n}$ denotes the range from the near-field target $k\in\mathcal{K}_{2}$ to the $n$-th antenna. Mathematically, $r_{k,n}$ can be expressed as
@@
for $k\in\mathcal{K}_{2}$, where $r_{k}$ represents the range between the target and the center of the BS. Further, by means of two-dimensional Fresnel approximation, $r_{k,n}$ in (5) can be approximated as [13]
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \!+\!\frac{n_{x}^{2}d_{0}^{2}(1-\alpha_{k}^{2})\!+\!n_{y}^{2}d_{0}^{2}(1-\beta_{k}^{2})}{2r_{k}}\!-\!\frac{n_{x}n_{y}d_{0}^{2}\alpha_{k}\beta_{k}}{r_{k}}.$
Then, the near-field channel steering vector for target $k\in\mathcal{K}_{2}$ can be rewritten as
@@
where $u_{n}(\alpha_{k},\beta_{k})\triangleq n_{x}d_{0}\alpha_{k}+n_{y}d_{0}\beta_{k}$ contains the angle information only, and $v_{n}(r_{k},\alpha_{k},\beta_{k})\triangleq\frac{n_{x}n_{y}d_{0}^{2}\alpha_{k}\beta_{k}}{r_{k}}-\frac{n_{x}^{2}d_{0}^{2}(1-\alpha_{k}^{2})+n_{y}^{2}d_{0}^{2}(1-\beta_{k}^{2})}{2r_{k}}$ is jointly determined by both the target angles and range.